OpenStudy (anonymous):

how do you solve 3/4 x =12

OpenStudy (anonymous):

first multiply 4 on both sides

OpenStudy (amistre64):

eloquently

OpenStudy (angela210793):

is x in the denom?

OpenStudy (anonymous):

u could divide each side by 3/4

OpenStudy (anonymous):

then divide 3 on both sides

OpenStudy (amistre64):

id try dividing it all by 12

OpenStudy (anonymous):

x=12(4/3)

OpenStudy (anonymous):

x=48/3

OpenStudy (anonymous):

x=<drumroll>16

OpenStudy (amistre64):

3x/4 = 12 /12 /12 ---------- x/16 = 1

jhonyy9 (jhonyy9):

3x =48 x=48/3 x=16